OpenStudy (anonymous):

Let f(x)=3x+5 and g(x)= -4x+7. Find (fog)(-4) The answers given are: A. 35 B. 74 C. -7 D. 23

OpenStudy (faiqraees):

Multiply f(x) by g(x) and then put x=4

OpenStudy (anonymous):

wait so I would do: 3x+5 * -4x+7?

OpenStudy (anonymous):

Input function g into function f, and then replace x for -4. (fog)(x)= -4(3x+5)+7 = -12x-20+7 = -12x-13 (fog)(-4) = -12(-4)-13 = 35

OpenStudy (anonymous):

Thank you so much! I have more questions like this and I'm seriously hating it right now lol. Thank you for explaining it to me!

OpenStudy (anonymous):

Happy to help! Good luck!

OpenStudy (anonymous):

What about for adding though??

OpenStudy (anonymous):

I have another one and it says: let f(x)= 4x-5 and g(x)= 3x+7. Find f(x)+g(x) and state its domain???

OpenStudy (anonymous):

Do you know what domain of a function means? I am guessing thats where the confusion is?

OpenStudy (anonymous):

That would mean all real numbers right?

OpenStudy (anonymous):

The confusion is more on how to write everything and then solve it. With these questions, there's an adding, dividing, and a multiplying one.

OpenStudy (anonymous):

I see, yes, this is just simple adding the functions. So, 4x-5+3x+7 = 7x+2. So, domain would be all real numbers. Correct

OpenStudy (anonymous):

OHH!! okay so for adding I would just add the both of them! Okay thank you!

OpenStudy (anonymous):

Yes, fog is different. Read fog as f as a function of g, just like f(x) is f as a function of x. Hope this clarifies.

OpenStudy (anonymous):

It definitely does. Thank you.